OpenStudy (anonymous):

CALCULUS HELP PLEASE I'm not sure what, but I'm doing something wrong on 2. A tomato is thrown upward from a bridge 25 m above the ground at 40 m/sec. (a) Give formulas for the acceleration, velocity, and height of the tomato at time t. (b) How high does the tomato go, and when does it reach its highest point? (c) How long is it in the air?

OpenStudy (anonymous):

Or, b, sorry.

OpenStudy (irishboy123):

first, show what you did/got for (a) and (b)....

OpenStudy (anonymous):

a) A=-9.8 V=-9.8t+40 h=-9.8t^2+40t+25 b)0=-9.8t+40 (I set it to zero velocity, or at max height) t=40/9.8 h=-9.8(40/9.8)^2+40(40/9.8)+25 h=-1600/9.8+1600/9.8+25 h=25? I'm messing up somewhere in there...

OpenStudy (amistre64):

h is wrong .... the first term

OpenStudy (anonymous):

OH, it is 4.9t^2 Thanks.

OpenStudy (anonymous):

Or, -4.9t^2

OpenStudy (amistre64):

-1/2 g t^2 yes

OpenStudy (anonymous):

Thank you very much!